# Topologically protected flatness in chiral moiré heterostructures

### v1

@@ V Implications for TMDs
While TBG has been the main focus up to now, let us expand on the implications of our results to twisted TMDs, which share the same chiral limit as TBG (Eq. 3) up to the addition of a mass term that describes the gap of the semiconducting monolayers [42]. Most of our analysis, particularly the decoupling into abelian components from Sec. III, goes through in the presence of this mass term. For twisted $\rm MoTe_{2}$ homobilayers, similar estimates as above show that the physics is effectively abelian for $\theta>2^{\circ}$, which lead to topologically protected flat bands at the first magic angles $\theta\sim 3.2^{\circ}$ [42]. Away from the chiral limit, the resulting flat band will acquire some dispersion due to additional orbital-preserving terms in the Hamiltonian but remain robust against disorder if the results of Fig. 7 can be generalized to twisted $\rm MoTe_{2}$ homobilayers. The mass term and the strong spin-orbit coupling of TMDs [36] strongly suppress all resonant non-chiral disorder – the only ones that spoil the protection of the flat bands, see Tab 1 – as they fully lift the degeneracy of the flat band to leave a single topologically protected valence flat band per spin flavor. Twisted TMDs, therefore, are ideal hosts for delicate correlated topological phases of matter. They host weakly dispersing Chern bands fully protected against all forms of disorder thanks to the combinations of the AS topological index theorem and their spectral gap.
@@ VI Conclusion
In this work, we have demonstrated that the flat band appearing at the first magic angle in twisted graphene and transition metal dichalcogenides bilayers, both in the chiral limit, can be understood as originating from Dirac particles evolving in an abelian gauge field. Their zero energy Landau levels become perfectly flat and topologically protected against disorder thanks to a topological index theorem when the moiré length and the effective magnetic length of the system match. This description is only possible when the inter-layer tunneling acting as a fictitious gauge field can be adiabatically decoupled into abelian components. Such decoupling occurs for large twist angles where the typical kinetic energy on the moiré scale is not negligible compared to the strength of tunneling and corresponds to $\theta>0.7^{\circ}$ for twisted bilayer graphene and $\theta>2^{\circ}$ for twisted $\rm MoTe_{2}$. The topological nature of the zeroth Landau level of chiral Dirac operators has critical consequences on the robustness of the magic angle against perturbations. In particular, we find that only the first magic angle benefits from the enhanced robustness to disorder granted by the Atiyah-Singer topological index. Therefore, we conclude that flat bands appearing at smaller twist angles are accidental, in the sense that they do not have an index and are strongly susceptible to spreading out in energy even for weak disorder on scales smaller than the moiré pattern. Let us finally highlight that all the flat bands at all magic angles are characterized by the same band topology (Chern number, Wannier obstructions, etc…) [65, 81], but only for those appearing at the first magic angle, the full Hamiltonian of TBG features an additional Atiyah-Singer topological index protecting the flat band.
@@ VII Acknowledgement
Appendix A Invertible transformation
In this appendix, we assume that a constant of motion $\vec{X}\cdot\vec{\mu}=X_{1}\mu_{1}+X_{2}\mu_{2}+X_{3}\mu_{3}$ solution of
$\left[\vec{X}\cdot\vec{\mu},\mathcal{D}\right]=0\quad\Longleftrightarrow\quad\partial_{\bar{z}}\vec{X}+\alpha\vec{X}\times\vec{\mathcal{A}}=0,$ (10)
exist in order to
  • show that $\vec{X}\cdot\vec{X}$ is uniform in space and real.
  • derive the form of the effective gauge potential $A_{z}^{\prime}$ appearing in Eq. 6.
  • prove that $\vec{X}$ cannot be carried by $\mu_{3}$ only away from high-symmetry points of the moiré unit cell.
A.1 Uniform and real $x^{2}$
We first demonstrate that $x^{2}=\vec{X}\cdot\vec{X}$ is uniform and real. We can first easily show that it is holomorphic
$\partial_{\bar{z}}(\vec{X}\cdot\vec{X})=2\vec{X}\cdot\partial_{\bar{z}}\vec{X}=-2\alpha\vec{X}\cdot(\vec{X}\times\mathcal{A})=0.$ (11)
Because $x$ is also additionally periodic (due to the periodicity of $\mathcal{A}$), it must be constant (uniform in space). Then, we use the $C_{2y}$ symmetry of the chiral Hamiltonian for TBG. This symmetry exchanges the layers and complex conjugates the gauge fields $A_{z}^{(j)}\to\left[A_{z}^{(j)}\right]^{*}$ for $j=1,2$, as can be derived from their explicit representation (Eq. 3). As a result, we find that $\vec{X}(-r_{1},r_{2})=\vec{X}^{*}(r_{1},r_{2})$, which ensures that all components of $\vec{X}$ are real on the $y$-axis where $r_{1}=0$ [we use $(r_{1},r_{2})$ for the position instead of $(x,y)$ to avoid confusion of the horizontal position with the $x^{2}=\vec{X}\cdot\vec{X}$], which results in a real $x^{2}$.
A.2 Abelianization
We then perform the calculation of the transformed Dirac operators $\mathcal{D}^{\prime}=U^{-1}\mathcal{D}U$ with
@@ A.2 Abelianization
A.3 Out-of-plane constant of motion
We finally prove by contradiction that $\vec{X}\cdot\vec{\mu}$ cannot be purely carried by $\mu_{3}$ at a point $r_{0}$ away from the origin and corners of the Wigner-Seitz moiré unit cell. We will rely on the fact that $x$ is constant and reason by contradiction. Assume that we can write $\vec{X}(r_{0})\cdot\vec{\mu}=x\mu_{3}$. Then, consider a small displacement $\eta=(\eta_{x},\eta_{y})$ away from this point. The equation defining the constant of motion Eq. 10 gives $X_{1}(r_{0}+\eta)=-x\alpha A_{z}^{(1)}(r_{0})\bar{\eta}$, $X_{2}(r_{0}+\eta)=x\alpha A_{z}^{(2)}(r_{0})\bar{\eta}$, and $X_{3}(r_{0}+\eta)=x$ with $\bar{\eta}=\eta_{x}-i\eta_{y}$. For the norm $x^{2}$ to be constant, we therefore need that $x^{2}=x^{2}[1+\alpha^{2}\bar{\eta}^{2}[(A_{z}^{(1)})^{2}(r_{0})+(A_{z}^{(2)})^{2}(r_{0}))]$, which requires $(A_{z}^{(1)})^{2}(r_{0})+(A_{z}^{(2)})^{2}(r_{0})=0$. This is only true for the field in Eq. 3 at the origin and corners of the Wigner-Seitz moiré unit cell, providing the desired contradiction.
Appendix B Numerical determination of magic angles
In this Appendix, we provide details on our numerical simulation of disorder in the chiral TBG Hamiltonian. We use the method introduced in Ref. [56], which we briefly summarize here for completeness.
@@ Appendix B Numerical determination of magic angles
Appendix C Small Momentum Perturbation
@@
Introducing harmonics that are periodic on an enlarged moiré unit cell in real space requires access to wave-vectors that are fractions of the original $\kappa_{1,n}$ appearing in the Bistritzer-MacDonald model (Eq. 3). To achieve this numerically, we rescale the momentum-space lattice of Fig. 8b by a factor $r=1/M$, with $M=3p+1$ and $p\in\mathbb{Z}$. This choice ensures that the reduced lattice still contains the $\kappa_{1,n}$. By convention, rescaling by a negative value $r$ corresponds to a rescaling by $|r|>0$ followed by a $180^{\circ}$ degree rotation, as illustrated in Fig. 9a for the case $r=-1/2$ (corresponding to $M=-2$, $p=-1$). The inter-layer harmonics $\hat{T}_{m}^{\prime}$ on the rescaled lattice with wave-vectors $\kappa_{m,n}^{\prime}$ can now be used to introduce disorder with correlation length larger than the moiré lattice constant. The definitions of $\hat{T}_{m}^{\prime}$ and $\kappa_{m,n}^{\prime}$ are the same as $\hat{T}_{m}$ and $\kappa_{m,n}$ (see App. B), albeit on the rescaled honeycomb lattice. We also highlight that $\kappa_{1,n}=\kappa_{|M|,n}^{\prime}$, which is also illustrated in Fig. 9a.
@@
where $\hat{T}^{\prime}_{|M|}$ is nothing but the terms appearing in the un-perturbed chiral Hamiltonian for TBG (Eq. 3), while $\hat{T}_{1}^{\prime}$ is the lowest accessible fractional harmonic on the rescaled lattice. The latter realizes a perturbation of momentum $q_{\rm pert}=|\kappa/M|$. In Fig. 9b, we show the derivative of the magic angles $|\partial\alpha/\partial\mu|$, numerically computed using a small disorder strength $\delta\mu=0.02$ (see Sec. IV.2), for the first three magic angles as a function of $M=\left|\frac{\kappa}{q_{\rm pert}}\right|=\left|\frac{\kappa}{\kappa_{1,0}^{\prime}}\right|$. We observe that, unlike the perturbation of larger momentum where only the first magic angle enjoyed protection, all magic angles respond similarly to disorder. Thus, there is no special protection of the first magic angle against perturbations of correlation length longer than the moiré lattice vector.

### v2

@@ V Implications for TMDs
Figure 6: Bandwidth of the lowest band of twisted TMDs modeled by a mass term $m=0.25$ for a) $w_{0}=0$ and b) $w_{0}=0.465w_{1}$ averaged over $50$ disorder realizations of strength $\delta\mu=0.1$ for the same distribution of $\mu_{m}$ as in Fig. 4, as a function of $\alpha$. The dashed line shows the same quantity in the absence of disorder. For b) we use the realistic parameter for WSe2 [44].
While TBG has been the main focus up to now, let us expand on the implications of our results to twisted TMDs, which share the same chiral limit as TBG (Eq. 3) up to the addition of a mass term that describes the gap of the semiconducting monolayers [44]. Because they are sublattice polarized and, therefore, eigenstates of the mass term, the flat bands of the chiral model at the magic angle remain completely flat under the introduction of the mass term but shift to non-zero positive/negative energy to describe the topmost valence and bottom-most conduction moiré bands of the system. Most of our analysis, particularly the decoupling into abelian components from Sec. III, goes through in the presence of this mass term. To show this, we calculate the averaged bandwidth of the lowest band of twisted TMDs with a mass term as a function of $\alpha$ in Fig. 6. One clearly sees that while the flatness of the first magic angle is robust, that of higher magic angles is destroyed by disorder. Away from the chiral limit, the resulting flat band will acquire some dispersion due to additional orbital-preserving terms in the Hamiltonian but remain robust against disorder if the results of Fig. 5 can be generalized to twisted $\rm MoTe_{2}$ homobilayers.
Our main result is that magic angles exists in twisted TMDs when their low-energy degrees of freedom are described by gapped Dirac cone with concentrated Berry curvature at the $K/K^{\prime}$ points, and that the corresponding flat bands carry a topological protection against disorder as long as the effective magnetic field obtained is well-behaved, which occurs in the abelian limit $\alpha\lesssim 1$ where the moiré tunneling in the original Hamiltonian lead to an almost abelian gauge field from the start (see discussion at the end of Sec. III.3). Our numerical results presented in Figs. 4 - 6 suggest that it is indeed the case. Taking $\alpha=1$ as the limit for topological protection, and using the effective Fermi velocity of graphene and $\rm MoTe_{2}$ [73], we expect the protection across all angles $\theta>0.7^{\circ}$ in TBG and $\theta>2^{\circ}$ for twisted $\rm MoTe_{2}$ homobilayers. In particular, the first magic angle of twisted $\rm MoTe_{2}$ at $\theta\sim 3.2^{\circ}$ is topologically protected [44].
The mass term and the strong spin-orbit coupling of TMDs [38] strongly suppress all resonant non-chiral disorder – the only ones that spoil the protection of the flat bands, see Tab 1 – as they fully lift the degeneracy of the flat band to leave a single topologically protected valence flat band per spin flavor. Twisted TMDs, therefore, are ideal hosts for delicate correlated topological phases of matter. They host weakly dispersing Chern bands fully protected against all forms of disorder thanks to the combinations of the AS topological index theorem and their spectral gap.
@@ VI Conclusion
In this work, we have demonstrated that the flat bands appearing at the magic angles in twisted graphene and transition metal dichalcogenides bilayers, both in the chiral limit, can be understood as originating from Dirac particles evolving in an abelian gauge field. Thanks to a topological index theorem, these zero-energy Dirac Landau levels become perfectly flat and topologically protected against disorder when the system’s moiré length and effective magnetic length match. Due to the pathological nature of the magnetic field at all but the first magic angle, we find that the topological protection offered by the index theorem can only be harnessed when the inter-layer tunneling is small enough compared to the typical Dirac kinetic energy at the moiré scale. We estimate the topologically protected region to span over all twist angles $\theta>0.7^{\circ}$ for twisted bilayer graphene and $\theta>2^{\circ}$ for twisted $\rm MoTe_{2}$. Only the first magic angle of both systems lies in this range, offering novel analytical insights into why delicate correlated phases such as fraction Chern insulators can be stabilized in these disordered systems.
@@ VII Acknowledgement
Appendix A Abelianization
We here assume that $\mathcal{D}$ admits a constant of motion of the form $\vec{X}\cdot\vec{\mu}$, i.e. $[\vec{X}\cdot\vec{\mu},\mathcal{D}]=0$, and derive calculate the specific form of the transformed Dirac operators $\mathcal{D}^{\prime}=U^{-1}\mathcal{D}U$ with
@@ Appendix A Abelianization
Appendix B Constant of motion
In this appendix, we show that the constant of motion $\vec{X}\cdot\vec{\mu}$ defined in Eq. 8 is traceless, of uniform norm, and reproduces Eq. 4.
In order to construct $\vec{X}\cdot\vec{\mu}$ as in Eq. 8, we need the existence of the two zero modes $|\psi_{\kappa}(\mathbf{r})\rangle$ and $|\psi_{\kappa^{\prime}}(\mathbf{r})\rangle$, which is guaranteed by $C_{2}T$ and $C_{3}$ symmtries [43]. If $C_{3}$ symmetry is broken in the system, there will still be at least two zero modes which can be used to construct the constant of motion, albeit not fixed at $\kappa$, $\kappa^{\prime}$ anymore. We assume the presence of $C_{3}$ throughout our derivation for simplicity.
We choose the relative phase between $|\psi_{\kappa}({r})\rangle$ and $|\psi_{\kappa^{\prime}}({r})\rangle$ such that $\operatorname{Tr}\vec{X}\cdot\vec{\mu}$ vanishes. This is because
$|\operatorname{Tr}(|\psi_{\kappa}({r})\rangle\langle\psi^{*}_{\kappa}(-{r})|)|=|\operatorname{Tr}(|\psi_{\kappa^{\prime}}({r})\rangle\langle\psi^{*}_{\kappa^{\prime}}(-{r})|)|=v_{F}(\alpha)$ (20)
is an integral of motion which is independent of $\mathbf{r}$ [43]. It only depends on $\alpha$ and corresponds to the Fermi velocity at the two Moire Dirac cones. Therefore, $\operatorname{Tr}(|\psi_{\kappa}({r})\rangle\langle\psi^{*}_{\kappa}(-{r})|)$ and $\operatorname{Tr}(|\psi_{\kappa^{\prime}}({r})\rangle\langle\psi^{*}_{\kappa^{\prime}}(-{r})|)$ can only be different up to a phase.
Let us calculate
$x^{2}(\alpha)=\frac{1}{2}\operatorname{Tr}(\vec{X}\cdot\vec{\mu})^{2}=v_{F}(\alpha)^{2}-\langle\psi^{*}_{\kappa}(-{r})|\psi_{\kappa^{\prime}}({r})\rangle\langle\psi^{*}_{\kappa^{\prime}}(-{r})|\psi_{\kappa}({r})\rangle=v_{F}(\alpha)^{2}$ (21)
In the third equity, we used the fact that the second term is an integral of motion which does not depend on $\mathbf{r}$ and it vanishes at $\mathbf{r}=\mathbf{r}_{0}=\frac{1}{3}(\mathbf{a}_{1}-\mathbf{a}_{2})$ so it is zero everywhere. Also, we have choose the overall phase of $\vec{X}\cdot\vec{\mu}$ such that $x^{2}$ is real.
More generally, we can construct $M$ with vanishing trace as
$M(\mathbf{r})=\begin{pmatrix}|\psi_{K}({r})\rangle&|\psi_{-K}({r})\rangle\end{pmatrix}R^{-1}\sigma_{z}R\begin{pmatrix}\langle\psi^{*}_{K}(-{r})|\\ \langle\psi^{*}_{-K}(-{r})\end{pmatrix}$ (22)
where $R$ represents any invertible $2\times 2$ matrix. It has the same property of $x=v_{F}$ and gives us an additional gauge degree of freedom. We choose $R=I_{2}$ (identity) without further notice. Also, when calculating the $U(1)$ gauge potential of Eq. 19, there is another degree of freedom to flip the overall sign of $\vec{X}$. In numerical calculations we find that $X_{3}$ is almost real. Therefore, we choose the sign such that $\operatorname{Re}X_{3}>0$. This helps to avoid singularity in the denominator $x+X_{3}$.
The wavefunction satisfies the translation property that $|\psi_{\kappa}\left({r}+{a}_{1,2}\right)\rangle=e^{i{\kappa}{a}_{1,2}}U_{\omega}|\psi_{\kappa}({r})\rangle$ with $U_{\omega}=\operatorname{diag}\left(1,\omega^{*}\right)$. Therefore, the constant of motion satisfies $\vec{X}\left({r}+{a}_{1,2}\right)\cdot\vec{\mu}=U_{\omega}\vec{X}({r})\cdot\vec{\mu}U^{\dagger}_{\omega}$. Also, because the symmetry property of the wavefunction: $C_{2y}T|\psi_{\kappa}\rangle=|\psi_{\kappa^{\prime}}\rangle$, we have $C_{2y}T\vec{X}\cdot\vec{\mu}C_{2y}T=-\vec{X}\cdot\vec{\mu}$. It is then straightforward to see that $\mathcal{D}^{\prime}({r}+{a}_{1,2})=U_{\omega}\mathcal{D}^{\prime}({r})U_{\omega}^{\dagger}=\mathcal{D}^{\prime}$, because of the fact that $\mathcal{D}^{\prime}$ is diagonal. This is indeed the translation property we expect for effective magnetic field with zero total flux in one unit cell. Also, one can check that $\mathcal{D}^{\prime}$ is odd under $C_{2y}T$, which leads to the property of the effective magnetic field that $B^{\prime}(-x,y)=-B^{\prime}(x,y)$.
Appendix C Momentum space spread and stability of the magic angles
Fig. 7 shows the wavefunction at the $\gamma$ point of the moiré Brillouin zone on the momentum-space lattice for the first three magic angles. It is evident that the momentum distribution of the wavefunction is more spread out for higher magic angles. This broader distribution correlates with the stability of magic angles against perturbations shown in Fig. 4. In the following, we present a perturbative argument to explain this behavior.
Figure 7: Wavefunction at the $\gamma$ point of the Moire BZ on the momentum space lattice for the first three magic angles.
The magic angles are determined by the eigenvalues of the operator $\Lambda$, which is expressed as
$\Lambda=\hat{D}^{-1}\hat{T}_{+}\hat{D}^{-1}\hat{T}_{-},\quad[\hat{D}]_{k,k^{\prime}}=\delta_{k,k^{\prime}}\frac{1}{k},\quad[\hat{T}_{+}]_{k,k^{\prime}}=\omega^{j}\delta_{k,k^{\prime}-q_{j}},\quad\hat{T}_{-}=\hat{T}_{+}^{\dagger}.$ (23)
Here, $\Lambda$ is a non-Hermitian operator with left and right eigenvectors that satisfy the relations
$\Lambda|\psi_{\alpha}\rangle=\frac{1}{\alpha^{2}}|\psi_{\alpha}\rangle,\quad\langle\tilde{\psi}_{\alpha}|\Lambda=\langle\tilde{\psi}_{\alpha}|\frac{1}{\alpha^{2}},\quad\langle\tilde{\psi}_{\alpha}|\psi_{\alpha^{\prime}}\rangle=\delta_{\alpha,\alpha^{\prime}}.$ (24)
The magic angles are the $\alpha$ obtained from this equation, which may be real or imaginary. Next, we introduce a perturbation of the form
$\hat{T}_{\pm}\rightarrow\hat{T}_{\pm}+\delta\hat{T}_{\pm},$ (25)
where $\delta\hat{T}_{\pm}$ couples momenta $k$ to $k+Q$, and $Q$ represents a higher momentum, i.e., $|Q|\geq|q_{j}|$. The first-order correction to the eigenvalues due to this perturbation is given by
$\langle\tilde{\psi}_{\alpha}|\hat{D}^{-1}\hat{T}_{+}\hat{D}^{-1}\delta\hat{T}_{-}|\psi_{\alpha}\rangle+\langle\tilde{\psi}_{\alpha}|\hat{D}^{-1}\delta\hat{T}_{+}\hat{D}^{-1}\hat{T}_{-}|\psi_{\alpha}\rangle,$ (26)
which corresponds to a scattering processes in the momentum space. The value of this expression depends sensitively on the spread of the wavefunction $|\psi_{\alpha}\rangle$. More precisely, for a wavefunction of the form $|\psi_{\alpha}\rangle=\sum_{G}v_{G}^{\alpha}|G\rangle$ or $\langle\psi_{\alpha}|=\sum_{G}w_{G}^{\alpha}\langle G|$, the correction has terms like
$\sum\limits_{G,G^{\prime}}w_{G^{\prime}}v_{G}\dfrac{1}{G^{\prime}}\left(\sum\limits_{j}\dfrac{\omega^{j}}{G^{\prime}+q_{j}}\right)\langle G^{\prime}+q_{j}|\delta T_{-}|G\rangle$ (27)
which clearly vanish if $\delta T_{-}$ connects $G$ to $G+Q$ where $|Q|\gg k_{\xi}$ where $k_{\xi}$ is the extent of the wavefunction in momentum space. As we show in Fig.7, the first magic angle, $\alpha=0.586$, has the most localized wavefunction with the smallest $k_{\xi}\approx|q_{j}|$. Consequently, the first-order correction vanishes for almost all higher harmonics, consistent with Fig. 4 in the main text. The same reasoning can further be applied to the second-order correction, which would scale as $1/Q^{2}$. The suppression arises from the two powers of the propagator $\hat{D}^{-1}$ in Eq. 26.
In contrast, the second and higher magic angles have wavefunctions that are more spread out in momentum space. There are many contributions to Eq. 26 and as a result, these wavefunctions are more susceptible to the effects of disorder from this perturbative standpoint.
Appendix D Details of disorder calculation
In this Appendix, we provide details on our numerical simulation of disorder in the chiral TBG Hamiltonian. We use the method introduced in Ref. [55], which we briefly summarize in Sec. D.3 for completeness. We then analyze the results of the main text showing the robustness of the first magic angle using localization in momentum space (Sec. C). Finally, we provide additional results for disorders defined on length scale smaller than the moiré lattice constant (Sec. D.3).
D.1 Numerical method
@@
D.2 Disorder $C_{3}$-symmetric on average
Figure 9: Average bandwidth of the lowest band $<W>$ versus $\alpha$ for a) the chiral model ($w_{0}=0$) and b) away from chiral limit $w_{0}=0.8w$ with the presence of $C_{3}$ breaking disorder. The black dashed line denotes the quantity without disorder. Disorder strength $\delta\mu=0.1$, $R_{k}=5\sqrt{3}|\kappa|$ and $50$ disorder realizations are averaged over.
The disorder we introduced in Sec. D.3 is $C_{3}$ symmetric by definition. The robustness of the first magic angle against these disorder has been elaborate in Sec. IV.2 in the main text. Here we investigate disorder that breaks $C_{3}$ symmetry, and provide numerical evidence that the first magic angle enjoys the same protection while higher magic angles do not. To implement $C_{3}$ breaking disorder, we simply modify Eq. (30) that $\delta\hat{T}=\sum_{mn}\mu_{mn}\hat{T}_{mn}=$ with $\hat{T}_{mn}=e^{i(\kappa_{m,n}\cdot r+\phi_{m,n})}$ and $\{\mu_{mn}\}$ is a random vector of small norm $\delta\mu\equiv|\mu|=[\sum_{mn}\mu_{mn}^{2}]^{1/2}\ll 1$ as before. With broken $C_{3}$ symmetry, the positions of the Dirac cones may be shifted so that it is difficult to characterize the Fermi velocity. Instead, we calculate the average bandwidth of the lowest band, as shown in Fig. 9. The calculation conveys a clear message that only the first magic angle is robust against higher harmonic disorder, even to those that breaks $C_{3}$ symmetry.
D.3 Small momentum perturbation
@@
Introducing harmonics that are periodic on an enlarged moiré unit cell in real space requires access to wave-vectors that are fractions of the original $\kappa_{1,n}$ appearing in the Bistritzer-MacDonald model (Eq. 3). To achieve this numerically, we rescale the momentum-space lattice of Fig. 8b by a factor $r=1/M$, with $M=3p+1$ and $p\in\mathbb{Z}$. This choice ensures that the reduced lattice still contains the $\kappa_{1,n}$. By convention, rescaling by a negative value $r$ corresponds to a rescaling by $|r|>0$ followed by a $180^{\circ}$ degree rotation, as illustrated in Fig. 10a for the case $r=-1/2$ (corresponding to $M=-2$, $p=-1$). The inter-layer harmonics $\hat{T}_{m}^{\prime}$ on the rescaled lattice with wave-vectors $\kappa_{m,n}^{\prime}$ can now be used to introduce disorder with correlation length larger than the moiré lattice constant. The definitions of $\hat{T}_{m}^{\prime}$ and $\kappa_{m,n}^{\prime}$ are the same as $\hat{T}_{m}$ and $\kappa_{m,n}$ (see App. D), albeit on the rescaled honeycomb lattice. We also highlight that $\kappa_{1,n}=\kappa_{|M|,n}^{\prime}$, which is also illustrated in Fig. 10a.
@@
where $\hat{T}^{\prime}_{|M|}$ is nothing but the terms appearing in the un-perturbed chiral Hamiltonian for TBG (Eq. 3), while $\hat{T}_{1}^{\prime}$ is the lowest accessible fractional harmonic on the rescaled lattice. The latter realizes a perturbation of momentum $q_{\rm pert}=|\kappa/M|$. In Fig. 10b, we show the derivative of the magic angles $|\partial\alpha/\partial\mu|$, numerically computed using a small disorder strength $\delta\mu=0.02$ (see Sec. IV.2), for the first three magic angles as a function of $M=\left|\kappa/q_{\rm pert}\right|=\left|\kappa/\kappa_{1,0}^{\prime}\right|$. We observe that, unlike the perturbation of larger momentum where only the first magic angle enjoyed protection, all magic angles respond similarly to disorder. We now understand this behavior from the momentum space picture presented in the previous section. There is no special protection of the first magic angle against perturbations of correlation length longer than the moiré lattice vector.